A bank loaned out $7,5000 part at 12% annual interest and the rest at 8% annual interest. The total interest earned for both loans was $680.00. How much was loaned at each rate?
Answer by mananth(16946)

A bank loaned out $7,500 part at 12% annual interest and the rest at 8% annual interest. The total interest earned for both loans was $680.00. How much was loaned at each rate?
12.00% per annum ------------- x
8.00% per annum ------------ y
total investment 75000
Interest----- $680.00
...

x+y=7500 ------------------------1
12.00% x+8.00% y=$680.00
Multiply by 100
12x+8y= $68,000.00 --------2

Multiply (1) by -12
we get

-12x-12 y=-90000

Add this to (2)

-4y=-$22,000.00

divide by-4

y=$5,500.00investment at 8.00%
Balance $2,000.00 investment at 12.00%